Please help me with this: Find F9x) at the indicated value of x.
f(x)=-6x - 1, x = -3
thank you

Answer by stanbon(75887)

Please help me with this: Find F9x) at the indicated value of x.
f(x)=-6x - 1, x = -3
---
f(-3) = -6*(-3) - 1
---
f(-3) = 18 - 1
---
f(-3) = 17
==============
Cheers,
Stan H.
============